OpenStudy (anonymous):

PLEASE HELP! Medal & Fan Given. Belinda wants to invest $1000. The table below shows the value of her investment under two different options for two different years:

OpenStudy (anonymous):

Number of years 1 2 3 Option 1 (amount in dollars) 1300 1600 1900 Option 2 (amount in dollars) 1120 1254.40 1404.93

OpenStudy (anonymous):

Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Part C: Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Which option should Belinda choose? Explain your answer, and show the investment value after 20 years for each option. (4 points)

OpenStudy (mokeira):

What is the question?

OpenStudy (anonymous):

i just put them up

OpenStudy (mokeira):

ooh..seen it sory

OpenStudy (anonymous):

it's cool

OpenStudy (puzzler7):

Part A: Linear for option 1, because it increases at a constant rate, exponential for option 2 because it is multiplied by a rate each time. Part B: Option 1: f(n) = 1000 +300n, because it starts at 1000 and increases 300 every year. Option 2: f(n) = 1000*1.12^n, because it gets multiplied by 1.12 every time. Part C: Option 1: f(20) = 1000+300*20 = 7000 Option 2: f(20) = 1000*1.12^20 = 9646.29 Option 2 is better.

OpenStudy (anonymous):

Thank you :)